A 2.0 L flask is filled with 0.30 mol SO3 , 0.40 mol of SO2 and 0.50 mol of O2 and allowed to reach equilibrium. Assume the temperature of the mixture is chosen so that Kc = 0.34. Predict the effect on the concentration of SO3 as the equilibrium is achieved by using Q, the reaction quotient.

2SO3(g) —> 2SO2(g) + O2(g)
Answer #1

2SO3(g) —> 2SO2(g) + O2(g)

[SO3] = 0.30 / 2 = 0.15 M

[SO2] = 0.40 / 2 = 0.20 M

[O2] = 0.50 / 2 = 0.25 M

reaction quotient Q = [SO2]^2[O2] / [SO3]^2

                                = (0.20)^2 x 0.25/ (0.15)^2

                                 = 0.44

Kc = 0.34

Qc > Kc

so the equilibrium shift to left .

concentration of SO3 increases.